# Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics

### v2

@@ Abstract
Parallel Data Curation (PDC) techniques aim to filter out noisy parallel sentences from web-mined corpora. Ranking sentence pairs using similarity scores on sentence embeddings derived from Pre-trained Multilingual Language Models (multiPLMs) is the most common PDC technique. However, previous research has shown that the choice of the multiPLM significantly impacts the quality of the filtered parallel corpus, and the NMT models trained using such data show a disparity across multiPLMs. This paper shows that this disparity is due to different multiPLMs being biased towards certain types of sentence pairs, which are treated as noise from an NMT point of view. We show that such noisy parallel sentences can be removed to a certain extent by employing a series of heuristics. The curated corpus leads to producing better results by NMT models while minimizing the disparity across multiPLMs. We publicly release the source code and the curated datasets [^1] Source codes and Datasets will be released upon acceptance.
@@
Aloka Fernando1, Nisansa de Silva1, Menan Velayuthan1, Charitha Rathnayake1 Surangika Ranathunga 2 {alokaf,NisansaDdS,velayuthan.22,charitha.18}@cse.mrt.ac.lk s.ranathunga@massey.ac.nz 1Dept. of Computer Science and Engineering, University of Moratuwa,10400, Sri Lanka 2Massey University, Palmerston North, New Zealand, 4443 Correspondence: alokaf@cse.mrt.ac.lk
@@ 1 Introduction
Parallel data mined from the web at scale is often considered an alternative to human-created data in training Neural Machine Translation (NMT) models Costa-jussà et al. (2022); Bañón et al. (2020). CCAligned El-Kishky et al. (2020), CCMatrix Schwenk et al. (2021) and ParaCrawl Bañón et al. (2020) are examples of such web-mined corpora, which cover Low-Resource Languages (LRLs) as well. However, quality audits of these corpora, as shown in Table 1, reveal substantial noise, which negatively impacts the NMT models Khayrallah and Koehn (2018).
Noise Category A B C D E
@@
Table 1: Noise types in parallel corpora, as identified by Khayrallah and Koehn (2018) (A), Bane et al. (2022) (B), Herold et al. (2022) (C), Kreutzer et al. (2022) (D) and Ranathunga et al. (2024) (E).
@@
Parallel Data Curation (PDC) aims at improving the quality of such web-mined parallel corpora by filtering high-quality parallel sentence pairs. The importance of PDC for LRLs has been emphasised with the introduction of WMT shared tasks Sloto et al. (2023); Koehn et al. (2020, 2019). Initiated by the work of  Chaudhary et al. (2019), recent PDC techniques follow a scoring and ranking mechanism using embeddings obtained from a Multilingual Language Model (multiPLM). Then, a subset of the top-ranked sentence pairs is selected to train the NMT model. However, the quality of these top-ranked pairs depends on the selected multiPLM (Ranathunga et al., 2024; Moon et al., 2023).
@@
To investigate this further, we obtain embeddings from three multiPLMs: LASER3 Heffernan et al. (2022), XLM-R Conneau et al. (2020), and LaBSE Feng et al. (2022), calculate the semantic similarity between each parallel sentence pair in the CCMatrix Schwenk et al. (2021) and CCAligned El-Kishky et al. (2020) datasets and rank them in descending order. Then we train NMT models (Section 4.4) using the top-ranked 100k sentence pairs from each corpus. Experiments are carried out for English-Sinhala (En-Si), English-Tamil (En-Ta) and Sinhala(Si-Ta). As shown in Figure 1, there is a significant disparity among the results, mainly for En-Si and En-Ta language pairs.
@@
Figure 1: Baseline NMT scores using top-ranked sentence pairs from CCMatrix and CCAligned, using embeddings obtained from LASER3, XLM-R, and LaBSE.
@@
A manual inspection of the En-Si top-ranked 1000 sentences reveal that different multiPLMs prioritise different sentence characteristics when ranking parallel sentences. For example, sentence pairs ranked top by XLM-R are mostly short and contain overlapping text such as numbers, acronyms, and URLs. LaBSE also favours sentences with numbers and date overlaps, while LASER3 prioratizes relatively better full length sentence-pairs. The impact of the nature of the top-ranked sentences is evident in Figure 1, where results related to LASER3 outperform its two counterparts by a significant margin, in most of the experiments.
@@
In order to carry out a more quantitative evaluation, we randomly selected 100 sentence pairs from the top 1000 sentences from the aforementioned ranked corpora and carried out a human evaluation. Ranathunga et al. (2024)’s is the most comprehensive error taxonomy available for this task. However, to better capture the noise identified during the manual observation, we extended this error taxonomy (Section 3). Human evaluation results in Table 2 show that compared to LASER3, the error category percentage (E) is excessively high for XLM-R and LaBSE. The most contributing types of errors are untranslated text (UN), short sentences (CS) and sentences with number/acronym/URL overlaps (CCN). Examples for these are shown in Table 10 in Appendix F.
@@
LASER3-Before 8.00% 27.00% 2.00%     37.00%     14.00% 14.00% 34.00% 1.00% 0.00% 0.00%     63.00%     17.00% 7.00% 4.00%     28.00%     7.00% 10.00% 55.00% 0.00% 0.00% 0.00%     72.00%     0.00% 3.00% 2.00%     5.00%     0.00% 0.00% 95.00% 0.00% 0.00% 0.00%     95.00%
LASER3-After 16.00% 56.00% 1.00%     73.00%     13.00% 4.00% 10.00% 0.00% 0.00% 0.00%     27.00%     39.00% 39.00% 7.00%     85.00%     0.00% 7.00% 8.00% 0.00% 0.00% 0.00%     15.00%     6.00% 61.00% 20.00%     87.00%     0.00% 3.00% 10.00% 0.00% 0.00% 0.00%     13.00%
XLM-R-Before 1.00% 10.00% 0.00%     11.00%     40.00% 19.00% 29.00% 0.00% 1.00% 0.00%     89.00%     1.00% 0.00% 0.00%     1.00%     13.00% 4.00% 80.00% 2.00% 0.00% 0.00%     99.00%     0.00% 0.00% 2.00%     2.00%     3.00% 5.00% 90.00% 0.00% 0.00% 0.00%     98.00%
XLM-R - After 0.00% 20.00% 2.00%     22.00%     13.00% 29.00% 35.00% 0.00% 1.00% 0.00%     78.00%     3.00% 8.00% 26.00%     37.00%     0.00% 2.00% 53.00% 8.00% 0.00% 0.00%     63.00%     0.00% 39.00% 31.00%     70.00%     1.00% 3.00% 21.00% 4.00% 0.00% 1.00%     30.00%
LaBSE - Before 4.00% 6.00% 0.00%     10.00%     74.00% 7.00% 9.00% 0.00% 0.00% 0.00%     90.00%     13.00% 2.00% 0.00%     15.00%     63.00% 14.00% 8.00% 0.00% 0.00% 0.00%     85.00%     0.00% 9.00% 2.00%     11.00%     34.00% 7.00% 48.00% 0.00% 0.00% 0.00%     89.00%
LaBSE - After 29.00% 33.00% 0.00%     62.00%     2.00% 32.00% 4.00% 0.00% 0.00% 0.00%     38.00%     87.00% 7.00% 3.00%     97.00%     0.00% 1.00% 2.00% 0.00% 0.00% 0.00%     3.00%     36.00% 53.00% 4.00%     93.00%     1.00% 3.00% 2.00% 1.00% 0.00% 0.00%     7.00%
@@
LASER3-Before 3.00% 24.00% 3.00%     30.00%     34.00% 19.00% 17.00% 0.00% 0.00% 0.00%     70.00%     2.00% 22.00% 8.00%     32.00%     13.00% 30.00% 23.00% 2.00% 0.00% 0.00%     68.00%     2.00% 23.00% 18.00%     43.00%     13.00% 27.00% 17.00% 0.00% 0.00% 0.00%     57.00%
LASER3-After 5.00% 79.00% 2.00%     86.00%     0.00% 9.00% 4.00% 1.00% 0.00% 0.00%     14.00%     13.00% 58.00% 14.00%     85.00%     0.00% 0.00% 13.00% 2.00% 0.00% 0.00%     15.00%     3.00% 67.00% 10.00%     80.00%     0.00% 8.00% 12.00% 0.00% 0.00% 0.00%     20.00%
XLM-R-Before 0.00% 0.00% 2.00%     2.00%     48.00% 49.00% 0.00% 0.00% 0.00% 1.00%     98.00%     2.00% 0.00% 0.00%     2.00%     72.00% 20.00% 6.00% 0.00% 0.00% 0.00%     98.00%     0.00% 8.00% 4.00%     12.00%     42.00% 16.00% 15.00% 8.00% 0.00% 7.00%     88.00%
XLM-R - After 20.00% 33.00% 4.00%     57.00%     1.00% 22.00% 19.00% 0.00% 1.00% 0.00%     43.00%     18.00% 18.00% 20.00%     56.00%     0.00% 6.00% 34.00% 4.00% 0.00% 0.00%     44.00%     6.00% 46.00% 30.00%     82.00%     0.00% 9.00% 9.00% 0.00% 0.00% 0.00%     18.00%
LaBSE - Before 0.00% 1.00% 0.00%     1.00%     69.00% 26.00% 3.00% 0.00% 0.00% 1.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 2.00% 0.00% 0.00% 0.00% 0.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 0.00% 0.00% 0.00% 0.00% 2.00%     99.00%
LaBSE - After 15.00% 34.00% 0.00%     49.00%     2.00% 43.00% 6.00% 0.00% 0.00% 0.00%     51.00%     45.00% 27.00% 3.00%     75.00%     1.00% 19.00% 5.00% 0.00% 0.00% 0.00%     25.00%     19.00% 45.00% 3.00%     67.00%     0.00% 22.00% 11.00% 0.00% 0.00% 0.00%     33.00%
Table 2: Human evaluation results for CCMatrix and CCAligned for En-Si, En-Ta and Si-Ta. Results are reported for LASER3, XLM-R, and LaBSE before and after applying heuristics. (C) - overall correct percentage considering $C$ (perfect translation), CN (near perfect) and CB (boilerplate). (E) - overall error percentage considering CCN (correct with overlaps), CS (correct but short sentence), X (wrong translation), UN (untranslated), WL (wrong language), NL (not a language).
We hypothesis that at least some of these noisy sentences can be filtered using rule-based heuristics. Although applying heuristics is a common approach to improving the quality of parallel corpora Sloto et al. (2023); Steingr´ımsson et al. (2023), the use of heuristics has not been consistent. They have limited their analysis to a single commonly used heuristic, or have varied the criteria or thresholds of the heuristics arbitrarily from one study to another. Further, their work does not explore the potential of combining multiple heuristics.
@@
  • We extend  Ranathunga et al. (2024)’s error taxonomy with a new error category to capture additional type of noise, which refers to sentence pairs with high-overlapping instances of numbers, acronyms, URLs.
  • We conduct a systematic study to analyse the impact of rule-based heuristics in filtering noisy sentences from the web-mined corpora and identify an optimal combination of heuristics that works across the corpora and languages considered in the study.
@@ 2.1 MultiPLMs for PDC
While employing a multiPLM for PDC is common, past research experimented with only one multiPLM at a time. For example, in the WMT2023 shared task, LASER2 was utilized to set the task baseline, whereas Steingr´ımsson (2023) and Gala et al. (2023) only used LaBSE. Therefore, the disparities across multiPLMs and biases specific to each multiPLM have not come into light.
On the other hand, studies conducted by Ranathunga et al. (2024) and Moon et al. (2023) reveal that using different multiPLMs for scoring and ranking parallel corpora, and training NMT models with the top-ranked corpora, results in a disparity. Moon et al. (2023) observe that this is due to biases in multiPLMs, which tend to rank noisy parallel sentences highly. However, there has been no study to identify these biases.
@@ 2.2 Identifying Noise in Web-mined Corpora
Recent research used categorical labels to annotate translation pairs, aiming at quantifying the noise types in web-mined corpora. Kreutzer et al. (2022) used their taxonomy to conduct manual audits on random samples from three web-mined datasets and reported substantial noise, specifically for LRLs. Ranathunga et al. (2024)’s taxonomy (See Table 9 in Appendix F) is an extension of Kreutzer et al. (2022)’s taxonomy. They first ranked the datasets based on embeddings from a multiPLM, and then selected random samples from top and bottom portions. Their human evaluation based on the taxonomy reported that the quality of the parallel sentences varies heavily depending on the selected portion.
@@ 2.3 Heuristic-based PDC
Commonly used rule-based heuristics can be categorised into four groups as described below:
De-duplication based (Dedup): Removing identical duplicates in the monolingual sides is a common practice Costa-jussà et al. (2022), while De-duplicating after removing non-alpha characters and punctuations Bala Das et al. (2023) could be found as its variants.
Length-based (sLength) : Gala et al. (2023) and Aulamo et al. (2023) have identified the removal of short sentences as a potential heuristic. Short sentences hinder NMT models in two ways Koehn and Knowles (2017): by providing insufficient syntactic and semantic information, or resulting in an overfitting situation.
LID-based (LID): Language Identification is used to remove fully/partially untranslated text and content in the wrong language Steingr´ımsson et al. (2023); Gala et al. (2023); Zhang et al. (2020).
Ratio-based : Ratio-based heuristics identify and remove sentences that show significant structural imbalances between the source and target sentences. It is based on the assumption that, well-aligned sentence pairs tend to maintain consistent ratios in terms of character count, word count, or token distribution. We observe three common types of ratio-based heuristics: (1) source-to-target sentence length ratio (STRatio(Rossenbach et al., 2018; Gale and Church, 1993), (2) alpha-only words to sentence words ratio (sentWRatio(Aulamo et al., 2020) and (3) alpha-only character ratio with respect to the sentence characters (sentCRatioHangya and Fraser (2018).
However, the impact of these heuristics in isolation and as a combination had not been evaluated systematically.
@@
CCN Perfect/near perfect translation where more than 30% of the overlapping content is numbers/acronyms/URLs/etc
@@
Table 3: A comparison of the improved taxonomy against Ranathunga et al. (2024)’s (only showing the changes).

### v3

@@ Abstract
Parallel Data Curation (PDC) techniques aim to filter out noisy parallel sentences from web-mined corpora. Ranking sentence pairs using similarity scores on sentence embeddings derived from Pre-trained Multilingual Language Models (multiPLMs) is the most common PDC technique. However, previous research has shown that the choice of the multiPLM significantly impacts the quality of the filtered parallel corpus, and the Neural Machine Translation (NMT) models trained using such data show a disparity across multiPLMs. This paper shows that this disparity is due to different multiPLMs being biased towards certain types of sentence pairs, which are treated as noise from an NMT point of view. We show that such noisy parallel sentences can be removed to a certain extent by employing a series of heuristics. The NMT models, trained using the curated corpus, lead to producing better results while minimizing the disparity across multiPLMs. We publicly release the source code and the curated datasets [^1] https://github.com/aloka-fernando/Heuristic_based_parallel_corpus_filtration.
@@
Aloka Fernando1, Nisansa de Silva1, Menan Velayuthan1, Charitha Rathnayake1 Surangika Ranathunga 2 {alokaf,NisansaDdS,velayuthan.22,charitha.18}@cse.mrt.ac.lk s.ranathunga@massey.ac.nz 1Dept. of Computer Science and Engineering, University of Moratuwa,10400, Sri Lanka 2School of Mathematical and Computational Sciences, Massey University, New Zealand Correspondence: alokaf@cse.mrt.ac.lk
@@ 1 Introduction
Parallel data mined from the web at scale is often considered an alternative to human-created data in training Neural Machine Translation (NMT) models Costa-jussà et al. (2022); Bañón et al. (2020). CCAligned El-Kishky et al. (2020), CCMatrix Schwenk et al. (2021) and ParaCrawl Bañón et al. (2020) are examples of such web-mined corpora, which cover Low-Resource Languages (LRLs) as well. As summarised in Table 1, quality audits of these corpora reveal different types of noise. Consequently, training NMT models on such noisy parallel data leads to low-quality translations Khayrallah and Koehn (2018).
Parallel Sentence Quality Category A B C D E
@@
Table 1: Parallel sentence quality categories used in quality audits by Khayrallah and Koehn (2018) (A), Bane et al. (2022) (B), Herold et al. (2022) (C), Kreutzer et al. (2022) (D) and Ranathunga et al. (2024) (E).
@@
Parallel Data Curation (PDC) aims at extracting high-quality parallel sentence pairs from noisy web-mined corpora. The importance of PDC for LRLs has been emphasised with the introduction of PDC shared tasks Sloto et al. (2023); Koehn et al. (2020, 2019). Initiated by the work of  Chaudhary et al. (2019), recent PDC techniques follow a scoring and ranking mechanism using embeddings obtained from a Multilingual Language Model (multiPLM). During the scoring step, the semantic similarity is calculated between the source and target sentence embeddings for each sentence pair. Then the sentence pairs are ranked in descending order of the similarity score. Finally, a subset of the top-ranked sentence pairs is selected to train the NMT model. However, the quality of these top-ranked pairs depends on the chosen multiPLM (Ranathunga et al., 2024; Moon et al., 2023).
@@
To investigate the impact of using different multiPLMs on the PDC task, we conduct an initial analysis. We obtain embeddings from three multiPLMs: LASER3 Heffernan et al. (2022), XLM-R Conneau et al. (2020), and LaBSE Feng et al. (2022), calculate the semantic similarity between each parallel sentence pair in the CCMatrix Schwenk et al. (2021) and CCAligned El-Kishky et al. (2020) datasets and rank them in descending order. Then we train NMT models (Section 4.4) using the top-ranked 100k sentence pairs from each corpus and report the ChrF++ scores. Experiments are carried out for English-Sinhala (En-Si), English-Tamil (En-Ta) and Sinhala-Tamil (Si-Ta) language-pairs. Sinhala, Tamil, and English belong to three distinct linguistic groups: Indo-Aryan, Dravidian, and Germanic (respectively). As shown in Figure 1, there is a significant difference, i.e. a disparity among these results, mainly for En-Si and En-Ta language pairs.
@@
Figure 1: Baseline NMT scores using top-ranked sentence pairs from CCMatrix and CCAligned corpora, with embeddings from LASER3, XLM-R, and LaBSE.
@@
A random inspection of the En-Si top-ranked sentences reveals that different multiPLMs prioritise different sentence characteristics when ranking parallel sentences. For example, sentence pairs ranked top with XLM-R embeddings are mostly short and have high overlapping text such as numbers, acronyms, and URLs. LaBSE embeddings also result in sentences with numbers and date overlaps, while LASER3 embeddings result in relatively better full-length sentence pairs. This observation sheds light on the disparity in NMT results shown in Figure 1 - the sentences ranked top when using XLM-R and LaBSE may have less linguistic content for NMT models to learn the translation.
@@
In order to carry out a more systematic evaluation, we randomly select 100 sentence pairs from the top 100K parallel sentences from the aforementioned ranked corpora and carried out a human evaluation. Ranathunga et al. (2024)’s is the most comprehensive taxonomy available for this task. To better capture the noise observed during human evaluation, we extend this taxonomy (Section 3). Human evaluation results (Table 2) indicate that the mean noise percentages are 98%, 95%, and 70% when the corpora are ranked using LaBSE, XLM-R, and LASER3 embeddings, respectively. Further, the noise types in the categories of untranslated text (UN), short sentences (CS) and sentences with high-overlapping non-translatable text (CCN) could be observed, contributing to the reported percentages. Examples for these noise types are shown in Table 10 in Appendix A. We provide a detailed discussion of these results in Section 5.3. These findings suggest that inherent biases in multiPLMs lead to noisy parallel sentences being ranked highly, which in turn contributes to disparities across NMT models.
@@
LASER3 BF 8% 27% 2%     37%     14% 14% 34% 1% 0% 0%     63%     17% 7% 4%     28%     7% 10% 55% 0% 0% 0%     72%     0% 3% 2%     5%     0% 0% 95% 0% 0% 0%     95%
AF 16% 56% 1%     73%     13% 4% 10% 0% 0% 0%     27%     39% 39% 7%     85%     0% 7% 8% 0% 0% 0%     15%     6% 61% 20%     87%     0% 3% 10% 0% 0% 0%     13%
XLM-R BF 1% 10% 0%     11%     40% 19% 29% 0% 1% 0%     89%     1% 0% 0%     1%     13% 4% 80% 2% 0% 0%     99%     0% 0% 2%     2%     3% 5% 90% 0% 0% 0%     98%
AF 0% 20% 2%     22%     13% 29% 35% 0% 1% 0%     78%     3% 8% 26%     37%     0% 2% 53% 8% 0% 0%     63%     0% 39% 31%     70%     1% 3% 21% 4% 0% 1%     30%
LaBSE BF 4% 6% 0%     10%     74% 7% 9% 0% 0% 0%     90%     13% 2% 0%     15%     63% 14% 8% 0% 0% 0%     85%     0% 9% 2%     11%     34% 7% 48% 0% 0% 0%     89%
AF 29% 33% 0%     62%     2% 32% 4% 0% 0% 0%     38%     87% 7% 3%     97%     0% 1% 2% 0% 0% 0%     3%     36% 53% 4%     93%     1% 3% 2% 1% 0% 0%     7%
@@
LASER3 BF 3% 24% 3%     30%     34% 19% 17% 0% 0% 0%     70%     2% 22% 8%     32%     13% 30% 23% 2% 0% 0%     68%     2% 23% 18%     43%     13% 27% 17% 0% 0% 0%     57%
AF 5% 79% 2%     86%     0% 9% 4% 1% 0% 0%     14%     13% 58% 14%     85%     0% 0% 13% 2% 0% 0%     15%     3% 67% 10%     80%     0% 8% 12% 0% 0% 0%     20%
XLM-R BF 0% 0% 2%     2%     48% 49% 0% 0% 0% 1%     98%     2% 0% 0%     2%     72% 20% 6% 0% 0% 0%     98%     0% 8% 4%     12%     42% 16% 15% 8% 0% 7%     88%
AF 20% 33% 4%     57%     1% 22% 19% 0% 1% 0%     43%     18% 18% 20%     56%     0% 6% 34% 4% 0% 0%     44%     6% 46% 30%     82%     0% 9% 9% 0% 0% 0%     18%
LaBSE BF 0% 1% 0%     1%     69% 26% 3% 0% 0% 1%     99%     0% 1% 0%     1%     97% 2% 0% 0% 0% 0%     99%     0% 1% 0%     1%     97% 0% 0% 0% 0% 2%     99%
AF 15% 34% 0%     49%     2% 43% 6% 0% 0% 0%     51%     45% 27% 3%     75%     1% 19% 5% 0% 0% 0%     25%     19% 45% 3%     67%     0% 22% 11% 0% 0% 0%     33%
Table 2: The Human evaluation results showing the average percentage for each annotation class for CCMatrix and CCAligned corpora for the En-Si, En-Ta and Si-Ta language pairs. The sample sentences have been obtained before (BF) and after (AF) applying the heuristics. (C) - overall correct percentage considering $C$ (perfect translation), CN (near perfect) and CB (boilerplate). (E) - overall error percentage considering CCN (Non-translatable overlaps), CS (correct but short sentence), X (wrong translation), UN (untranslated), WL (wrong language), NL (not a language).
We hypothesise that some of these noisy sentences can be filtered using rule-based heuristics. Although applying heuristics is a common approach to improving the quality of parallel corpora Sloto et al. (2023); Steingrímsson et al. (2023), the use of heuristics has not been consistent in the PDC tasks Steingrimsson (2023); Bala Das et al. (2023); Aulamo et al. (2020). Previous research either applied a single heuristic or a subset of commonly used heuristics during pre-processing, with threshold choices varying across studies. Further, they did not systematically analyse the impact of heuristic-based filtering on the NMT performance.
@@
  • We extend  Ranathunga et al. (2024)’s parallel sentence categorization taxonomy with a new error category to capture an additional type of noise, which refers to sentence pairs with high-overlapping non-translatable text such as numbers, acronyms, URLs, etc.
@@
  • We conduct a systematic study to analyse the impact of rule-based heuristics in filtering noisy sentences from the web-mined corpora and identify an optimal combination of heuristics that works across corpora and languages considered in the study.
@@ 2.1 MultiPLMs for PDC
While employing a multiPLM for PDC is common, existing research experimented with only one multiPLM at a time. For example, in the WMT2023 shared task Sloto et al. (2023), LASER2 was utilised to set the task baseline, whereas for the same task, Steingrimsson (2023) used LaBSE. Gala et al. (2023) also used LaBSE for their work. Therefore, the disparities across multiPLMs and biases specific to each multiPLM have not come to light.
On the other hand, studies conducted by Ranathunga et al. (2024) and Moon et al. (2023) reveal that using different multiPLMs for scoring and ranking parallel corpora, and training NMT models with the top-ranked corpora, results in a disparity. Moon et al. (2023) observe that this is due to biases in multiPLMs, which tend to rank noisy parallel sentences highly. However, there has been no systematic study to identify these biases.
@@ 2.2 Identifying Noise in Web-mined Corpora
Recent research used categorical labels to annotate translation pairs, aiming at quantifying the noise types in web-mined corpora. Kreutzer et al. (2022) used their taxonomy to conduct manual audits on random samples from three web-mined datasets and reported substantial noise, specifically for LRLs. Ranathunga et al. (2024)’s taxonomy (See Table 9 in Appendix A) is an extension of Kreutzer et al. (2022)’s taxonomy. They first ranked the datasets based on embeddings from a multiPLM, and then selected random samples from the top and bottom portions and conducted a quality audit. Their human evaluation reported that the quality of the parallel sentences varies heavily depending on the selected portion. These studies primarily focus on identifying general noise types; however, their effectiveness in quantifying the specific noise types to which multiPLMs are biased has not yet been evaluated to the best of our knowledge.
@@ 2.3 Heuristic-based PDC
In existing work, the commonly used rule-based heuristics can be categorised into four groups as described below:
Deduplication-based (Dedup): Removing identical duplicates from the monolingual sides is a common practice Costa-jussà et al. (2022). Additionally deduplicating after removing non-alpha characters and punctuations Bala Das et al. (2023) could be found as its variants. While this step has been applied during the pre-processing stage, an empirical study has not been conducted to evaluate its impact on the final NMT performance.
Length-based (sLength) : Gala et al. (2023) and Aulamo et al. (2023) have removed short sentences as a potential heuristic. Short sentences hinder NMT models in two ways Koehn and Knowles (2017): by providing insufficient syntactic and semantic information, and can result in an overfitting situation.
LID-based (LID): Language Identification is used to remove fully/partially untranslated text and content in the wrong language Steingrímsson et al. (2023); Gala et al. (2023); Zhang et al. (2020).
Ratio-based : Ratio-based heuristics identify and remove sentences that show significant structural imbalances between the source and target sentences. It is based on the assumption that, well-aligned sentence pairs tend to maintain consistent ratios in terms of character count, word count, or token distribution. We observe three common types of ratio-based heuristics: (1) source-to-target sentence length ratio (STRatio(Rossenbach et al., 2018; Gale and Church, 1993), (2) alpha-only words to sentence words ratio (sentWRatio(Velayuthan et al., 2024; Aulamo et al., 2020) and (3) alpha-only character ratio with respect to the sentence characters (sentCRatioHangya and Fraser (2018).
However, the impact of these heuristics in isolation and as a combination had not been evaluated systematically in the context of NMT.
@@
- CCN Perfect/near perfect translation where more than 30% of the overlapping content is non-translatable such as numbers/acronyms/URLs/email etc.
@@
Table 3: A comparison of the improved taxonomy against Ranathunga et al. (2024)’s (only showing the changes). See Table 9 in Appendix A for the full taxonomy.